# Gravitational $D$-Form Factor: The $\sigma$-Meson as a Dilaton confronted with Lattice Data

### v1

@@ 1 Introduction
Gravitational form factors probe the energy-momentum tensor for physical states, such as nucleons, through matrix elements of the form $\langle N(p^{\prime})|T_{\mu\nu}|N(p)\rangle$. The momentum transfer $q=p^{\prime}-p$ reveals the energy-momentum distribution of the nucleon just like electromagnetic form factors test the charge distribution. Defined a long time ago [1, 2], gravitational form factors have seen a growing interest (reviewed in [3, 4, 5, 6]). In part related to their experimental accessibility, through the first moment of generalised parton distributions [7], in deeply virtual Compton scattering [8, 9], near-threshold $J/\psi$ photoproduction [10], and $\gamma^{*}\gamma\to\pi\pi$ [11, 12, 13], among others. This has triggered lattice QCD investigations [14, 15, 16, 17, 18, 19] which provide the basis of this paper. In addition to experiment and lattice results, there are perturbative approaches in chiral theories at low $q^{2}$ [20, 21] and light-cone sum rules at high $q^{2}$ [22, 23, 24, 25, 26], dispersive analyses [27, 28, 29, 30], Skyrme-based models [31, 32, 33], light-front quark models [34, 35], and holographic models [36, 37, 38, 39].
@@
Conserved-current form factors often obey physical constraints at zero momentum transfer. An example is the charged-pion electromagnetic form factor $F_{+}(0)=1$, expressing charge conservation. The infrared interpretation of the gravitational $D$-form factor, associated with the internal pressure distribution [40], has remained elusive and puzzled the community for a long time [3, 6]. Our work offers an explanation in terms of the dilaton.
@@
The idea that strong interactions are governed by an infrared fixed point [41, 42, 43] has recently been reexamined in QCD, both in low-energy processes [44, 45] and more formally [46, 47, 48], by matching scaling dimensions of the underlying theory to those of the effective theory. Consistency regarding the quark-mass anomalous dimension with lattice simulations [49, 50, 51, 52, 53, 54, 55] (or fits thereto [56, 57]), phenomenological models, lower dimensional models [58] and ${\cal N}=1$ supersymmetric gauge theories [59, 47, 60] are established. The underlying idea of this scenario is that spontaneous scale symmetry breaking generates hadron masses, with the resulting Goldstone bosons, the dilaton, realising the corresponding Ward identities. [^1] Gauge theories dilatons are of interest elsewhere, e.g. the Higgs as a dilaton [61, 62, 63] or nuclear physics [64, 65, 66].
@@
The gravitational form factors, for a scalar $\varphi$, provides the ideal setting to illustrate these ideas
@@
with momenta $2{\cal P}\equiv p+p^{\prime}$ and $q$ as defined above. Energy conservation implies the model-independent constraint $A(0)=1$ and if one assumes that $D(q^{2})$ is regular for $q\to 0$, one recovers the standard textbook formula $\Theta(0)\equiv\Theta^{\rho}_{\phantom{\rho}\rho}(0)=2m_{\varphi}^{2}$. However, in the presence of a massless dilaton arising from spontaneous scale symmetry breaking, the dilaton pole
@@
modifies the textbook formula to satisfy the infrared conformal Ward identity $T^{\rho}_{\;\;\rho}=0$”, analogous to the Goldberger-Treiman mechanism, where the pion restores the chiral Ward identity [67]. [^2] In the axial singlet channel, the $\eta^{\prime}$ plays a similar role in QCD processes, see for instance [68]. For the dilaton it was proposed in ref. [69], and verified using the LSZ formalism  [70] and the effective theory  [71]. Consequently, one finds
@@
In this work, we test infrared conformality in a direct and physical way, independent of $\beta$-functions. Our strategy is to fit lattice data for the nucleon and pion gravitational $D$-form factor [16, 17] to the predictions of leading-order (LO) dilaton effective theory [71], which incorporates the dilaton Goldberger-Treiman mechanism described above.
The paper is organised as follows. In section 2 the gravitational form factors are defined for the nucleon and pion, and the LO dilaton predictions are given. Section 3 motivates a fit ansatz suitable in the Euclidean regime. In section 4 we present our fits, test the dilaton hypothesis and comment on the $D$-term. Conclusions and discussions follow in section 5. Appendix A explores the linear $\sigma$-model to clarify the relation between complex-valued and effective Euclidean poles, while appendix B sketches a multipole expansion in momentum space to motivate the fit ansatz.
@@ 2.1 Definition of nucleon and pion form factors
The gravitational form factors of the nucleon $N$ and the pion $\pi$ are defined, in the now-standard conventions, as
$\langle N(p^{\prime},s^{\prime})|T_{\mu\nu}|N(p,s)\rangle$ $\;=\;$ $\frac{1}{2m_{N}}\,\bar{u}(p^{\prime},s^{\prime})\Big{(}a_{\mu\nu}\,A^{N}(q^{2})+j_{\mu\nu}\,J^{N}(q^{2})+d_{\mu\nu}\,D^{N}(q^{2})\Big{)}u(p,s)\;,$
$\langle\pi(p^{\prime})|T_{\mu\nu}|\pi(p)\rangle$ $\;=\;$ $\qquad\qquad\qquad\big{(}a_{\mu\nu}\,A^{\pi}(q^{2})+d_{\mu\nu}\,D^{\pi}(q^{2})\big{)}\;,$ (2.1)
where $u(p,s)$ denotes the Dirac spinor, with ($\sigma_{\mu\nu}=\tfrac{i}{2}[\gamma_{\mu},\gamma_{\nu}]$ and $\sigma_{q\nu}=\sigma_{\mu\nu}q^{\mu}$)
@@
Lorentz structures ensuring translational invariance $q^{\mu}\Theta_{\mu\nu}(q)=0$. The normalisation $A(0)=1$ holds model-independently since $T_{\mu\nu}$ is the associated Noether current with conserved momenta $P_{\mu}=\int d^{3}xT_{\mu 0}$. Furthermore, for the nucleon one has $J^{N}(0)=\tfrac{1}{2}$, which reflects the absence of an anomalous gravitational magnetic moment, a consequence of the universality of gravity. [^3] The form factors $A^{\pi,N}$ and $J^{N}$ are the analogues of the electromagnetic pion, Dirac and Pauli form factors.
@@ 2.2 Leading order dilaton effective theory and $D(q^{2})$
The form factors have been evaluated in the LO dilaton effective theory in an arbitrary spacetime dimension $d$, including chiral corrections, in ref. [71]. Here, we consider it instructive to highlight the main effect, the dilaton pole in the $D$-form factor, using the nucleon as an example. Denoting the dilaton by $\sigma$ and the coset field by $\hat{\chi}=e^{-\sigma/F_{\sigma}}$, the relevant Lagrangian reads
@@
where $F_{\sigma}$, $R$, $\omega_{N}$ and $\Delta_{\mu}$ stand for the dilaton decay constant, the Ricci scalar, the nucleon Weyl weight (conformal charge) and the Weyl-covariant derivative, respectively. For the latter we refer the reader to original refs. [72, 41, 73], and [71] for an emphasis on generic Weyl-weight. We first focus on the improvement term proportional to the Ricci scalar. While irrelevant for scattering in flat space, it contributes to the energy-momentum tensor
@@
$D^{N}(q^{2})=\frac{r_{\sigma}^{N}}{q^{2}}\;,\quad r_{\sigma}^{N}=\frac{2}{3}\bar{u}(p)u(p)\,F_{\sigma}\,g_{\sigma NN}=\frac{4}{3}m_{N}^{2}\;,$ (2.7)
@@
The full LO computation, including linear quark mass effects, reads [71]
$\quad{D}^{N}(q^{2})=\frac{4}{3}\frac{\bar{m}_{N}^{2}}{q^{2}-m_{\sigma}^{2}}\;,\quad{D}^{\pi}(q^{2})=\frac{2}{3}\frac{q^{2}}{q^{2}-m_{\sigma}^{2}}-1\;,$ (2.8)
where $\bar{m}_{N}$ denotes the nucleon mass in the chiral limit. The physical nucleon mass can then be written as
@@
with $\Delta{m}_{N}$ the non-vanishing part due to $m_{u,d,s}\neq 0$. In the nucleon case the linear quark corrections vanish since the residue is of the form $m_{N}(\bar{m}_{N}-\Delta{m}_{N})=\bar{m}_{N}^{2}-(\Delta{m}_{N})^{2}$. Hence, further contributions are of ${\cal O}(m_{q}^{3/2})$ as in ordinary nucleon chiral perturbation theory [74]. The corrections in the pion case are ${\cal O}(m_{\pi}^{4}\ln m_{\pi})$ as usual. In the pion case it is worthwhile to point out that there is a soft-pion theorem constraining the LO trace to be [75, 71]
@@
Together with the model-independent normalisation $A^{\pi}(0)=1$, this implies the constraints444 Alternatively, one may write ${D}^{\pi}(q^{2})|_{\text{LO}}=\frac{2}{3}\frac{m_{\sigma}^{2}}{q^{2}-m_{\sigma}^{2}}-\frac{1}{3}$, more natural from the dispersive viewpoint.
@@
subject to higher order chiral corrections. We emphasise that, for nonzero pion mass, the constraint holds independently of any fixed-point interpretation and is therefore directly relevant for the present work. In fact, Donoghue and Leutwyler [75] employed the soft theorem together with the $\sigma$-meson dominance assumption to deduce this form. Consequently, the constraint $D^{\pi}(0)=-1$ provides a valuable handle for fixing the fit ansatz of the pion. In contrast, for the nucleon no analogous constraint exists, since constant or $q^{2}$-dependent shifts may arise from higher resonances or multi-particle contributions to the spectrum.
@@ 3 The $\sigma$-pole in the Euclidean
Describing the $\sigma$-pole is a non-trivial task since the $\sigma$-meson is perhaps the most complicated and mysterious resonance of QCD [76] and its pole is deep in the complex plane on the second sheet [77]
@@
Our main point is that the details of the $\sigma$-meson in the Minkowski domain are irrelevant in the deep Euclidean region, where a simple effective pole parametrisation proves sufficient. In section 3.2, we will revisit the issue and explain why alternative resonance parameterisations are not well suited.
@@ 3.1 The effect of a resonance in the deep Euclidean
where $\cal O$} denotes the operator, transitioning from $H(p)$ to $H(p^{\prime})$, and coupling to a stable particle $\varphi$ in the $q^{2}=(p-p^{\prime})^{2}$ channel. If we assume that $G(q^{2})$ is real for $q^{2}<m_{\varphi}^{2}$ (real analyticity), then it admits a dispersion representation of the form
$G(q^{2})=\int_{0}^{\infty}\frac{ds\,\rho_{G}(s)}{s-q^{2}-i0}=\frac{r_{0,\varphi}}{q^{2}-m^{2}_{\varphi}}+\delta G(q^{2})\;,$ (3.4)
@@
where $\rho_{G}(s)=\tfrac{1}{\pi}\textrm{Im}G(s)$ and $\delta G(s)$ encodes the remaining spectral contributions. In this case, the effect of the pole across the complex plane is simple and fully controlled by the residue $r_{0,\varphi}$. Now suppose we turn on a parameter that renders the particle $\varphi$ unstable. The pole then moves to the second sheet, $\sqrt{s_{\varphi}}=m_{\varphi}-\frac{i}{2}\Gamma_{\varphi}$, and one may define a complex-valued residue $r_{\varphi}$ through
$G^{(\text{II})}(q^{2})|_{q^{2}\approx s_{\varphi}}=\frac{r_{\varphi}}{q^{2}-s_{\varphi}}+{\cal O}(1)\;,$ (3.5)
where the superscript indicates the second-sheet continuation. While this definition is interesting in its own right, it is not the relevant quantity for analyzing the pole contribution in the deep Euclidean region, since there the form factor remains real.
Intuitively, we expect
$G(q^{2})|_{q^{2}\ll m_{\varphi}^{2}}\approx\frac{r_{\mathrm{E},\varphi}}{q^{2}-m_{\mathrm{E},\varphi}^{2}}+\delta\tilde{G}(q^{2})\;,$ (3.6)
where $r_{\mathrm{E},\varphi}$ and $m_{\mathrm{E},\varphi}$ denote effective Euclidean residue and mass, respectively. In particular, we anticipate
@@
$(i)\;\;\text{arg}(\sqrt{s_{\sigma}})\leftrightarrow\text{arg}(r_{\varphi})\;,\qquad(ii)\;\;r_{\mathrm{E},\varphi}\neq|r_{\varphi}|\;,$ (3.7)
that the phase in the pole correlates with the residue phase, and that the Euclidean residue can differ significantly from the modulus of the complex residue. More generally, one would expect $m_{\mathrm{E},\varphi}\approx|s_{\varphi}|$ and that $|r_{\mathrm{E},\varphi}/r_{\varphi}|<1$. Both (i) and (ii) hold true in the linear $\sigma$-model , presented in appendix A, while the more qualitative features remain valid across a wide range of parameters. In addition, a more formal treatment of this idea, in the spirit of a multipole expansion in momentum space, is given in appendix B.

### v2

@@ 1 Introduction
Gravitational form factors probe the energy-momentum tensor for physical states, such as nucleons, through matrix elements of the form $\langle N(p^{\prime})|T_{\mu\nu}|N(p)\rangle$. The momentum transfer $q=p^{\prime}-p$ reveals the energy-momentum distribution of the nucleon just like electromagnetic form factors test the charge distribution. Defined a long time ago Kobzarev:1962wt ; Pagels:1966zza , gravitational form factors have seen a growing interest (reviewed in Polyakov:2018zvc ; Burkert:2023wzr ; Lorce:2024ipy ; Lorce:2025oot ). In part related to their experimental accessibility, through the first moment of generalised parton distributions Ji:1996ek , in deeply virtual Compton scattering Burkert:2018bqq ; Duran:2022xag ; Goharipour:2025lep , near-threshold $J/\psi$ photoproduction GlueX:2023pev , and $\gamma^{*}\gamma\to\pi\pi$ Diehl:1998dk ; Belle:2015oin ; Kumano:2019vlv , among others. This has triggered lattice QCD investigations Shanahan:2018pib ; Pefkou:2021fni ; Hackett:2023rif ; Hackett:2023nkr ; Wang:2024lrm ; Abbott:2025irb which provide the basis of this paper. In addition to experimental and lattice results, there are perturbative approaches in chiral theories at low $q^{2}$ Belitsky:2002jp ; Alharazin:2023zzc and light-cone sum rules at high $q^{2}$ Anikin:2019kwi ; Tong:2021ctu ; Tong:2022zax ; Ozdem:2022zig ; Dehghan:2025ncw , dispersive analyses Broniowski:2024oyk ; Cao:2024zlf ; Broniowski:2025ctl ; Cao:2025dkv , Skyrme-based models Cebulla:2007ei ; Jung:2013bya ; Tanaka:2025pny ; Fukushima:2025jah , light-front quark models Chakrabarti:2020kdc ; Choi:2025rto , and holographic models Mamo:2021krl ; Mamo:2022eui ; Fujita:2022jus ; Sugimoto:2025btn ; Tanaka:2025znc .
@@
Conserved-current form factors often obey physical constraints at zero momentum transfer. An example is the charged-pion electromagnetic form factor $F_{+}(0)=1$, expressing charge conservation. The infrared interpretation of the gravitational $D$-form factor, associated with the internal pressure distribution Polyakov:2002yz , has remained elusive and puzzled the community for a long time Polyakov:2018zvc , as debated in Ji:2025gsq ; Ji:2025qax ; Lorce:2025oot . Our work offers an explanation in terms of the dilaton.
@@
The idea that strong interactions are governed by an infrared fixed point Isham:1970gz ; Zumino:1970tu ; Ellis:1971sa has recently been reexamined in QCD, both in low-energy processes Crewther:2013vea ; Crewther:2015dpa and more formally Zwicky:2023bzk ; Zwicky:2023krx ; Zwicky:2025moo , by matching scaling dimensions of the underlying theory to those of the effective theory. Consistency of the quark-mass anomalous dimension with lattice simulations DelDebbio:2015byq ; Hasenfratz:2020ess ; Fodor:2017gtj ; LatticeStrongDynamics:2018hun ; LatticeStrongDynamics:2023bqp ; Bennett:2024tex ; LatKMI:2025kti (or fits thereto Appelquist:2017wcg ; Appelquist:2017vyy ), phenomenological models, lower dimensional models Cresswell-Hogg:2025kvr and ${\cal N}=1$ supersymmetric gauge theories Terning:2006bq ; Zwicky:2023krx ; Shifman:2023jqn are established. The underlying idea of this scenario is that spontaneous scale symmetry breaking generates hadron masses, with the resulting Goldstone bosons, the dilaton, realising the corresponding Ward identities. [^1] Gauge theory dilatons are of interest elsewhere: the Higgs as a dilaton Matsuzaki:2012xx ; Dietrich:2005jn ; Cata:2018wzl ; Zwicky:2023krx or nuclear physics in dense matter Brown:1991kk ; Rho:2021zwm ; Rho:2024jdo .
@@
The gravitational form factors for a scalar $\varphi$ provide the ideal setting to illustrate these ideas
@@
with momenta $2{\cal P}\equiv p+p^{\prime}$ and $q$ as defined above. Energy conservation implies the model-independent constraint $A(0)=1$ and if one assumes that $D(q^{2})$ is regular for $q\to 0$, one recovers the standard textbook formula $\Theta(0)\equiv\Theta^{\rho}_{\phantom{\rho}\rho}(0)=2m_{\varphi}^{2}$ Donoghue:1992dd (for the normalisation $\langle\varphi(p^{\prime})|\varphi(p)\rangle=2E_{p}(2\pi)^{3}\delta(\vec{p}-\vec{p}\,^{\prime})$). However, in the presence of a massless dilaton arising from spontaneous scale symmetry breaking, the dilaton pole
@@
modifies the textbook formula to satisfy the infrared conformal Ward identity $T^{\rho}_{\;\;\rho}=0$”, analogous to the Goldberger-Treiman mechanism, where the pion restores the chiral Ward identity Donoghue:1992dd . [^2] In the axial singlet channel, the $\eta^{\prime}$ plays a similar role in QCD processes, see for instance Tarasov:2020cwl ; Tarasov:2021yll ; Tarasov:2025mvn ; Bhattacharya:2022xxw ; Bhattacharya:2023wvy . For the dilaton it was proposed in ref. Gell-Mann:1969rik , and verified using the LSZ formalism  DelDebbio:2021xwu and the effective theory  Zwicky:2023fay . Consequently, one finds
@@
In this work, we test infrared conformality in a direct and physical way, independent of $\beta$-functions. Our strategy is to fit lattice data for the nucleon and pion gravitational $D$-form factor Hackett:2023rif ; Hackett:2023nkr to the predictions of leading-order (LO) dilaton effective theory Zwicky:2023fay , which incorporates the dilaton Goldberger-Treiman mechanism described above.
The paper is organised as follows. In section 2 the gravitational form factors are defined for the nucleon and pion, and the LO dilaton predictions are given. Section 3 motivates a fit ansatz suitable in the Euclidean regime. In section 4 we present our fits, test the dilaton hypothesis and comment on the $D$-term. Conclusions and discussions follow in section 5. Appendix A explores the linear $\sigma$-model to clarify the relation between complex-valued and effective Euclidean poles, while appendix B sketches a multipole expansion in momentum space to motivate the fit ansatz. In appendix C we present additional plots putting our fits into perspective with other work.
@@ 2.1 Definition of nucleon and pion form factors
The gravitational form factors of the nucleon $N$ and the pion $\pi$ are parametrised by
$\Theta^{N}_{\mu\nu}(q)$ $\;=\;$ $a_{\mu\nu}\,A^{N}(q^{2})+j_{\mu\nu}\,J^{N}(q^{2})+d_{\mu\nu}\,D^{N}(q^{2})\;,$
$\Theta^{\pi}_{\mu\nu}(q)$ $\;=\;$ $a_{\mu\nu}\,A^{\pi}(q^{2})+d_{\mu\nu}\,D^{\pi}(q^{2})\;,$ (2.1)
where the left hand side is defined by
$\frac{1}{2m_{N}}\,\bar{u}(p^{\prime},s^{\prime})\Theta^{N}_{\mu\nu}(q)u(p,s)\equiv\langle N(p^{\prime},s^{\prime})|T_{\mu\nu}|N(p,s)\rangle\;,\quad\Theta^{\pi}_{\mu\nu}(q)\equiv\langle\pi(p^{\prime})|T_{\mu\nu}|\pi(p)\rangle\;,$ (2.2)
and $u(p,s)$ denotes the Dirac spinor, with ($\sigma_{\mu\nu}=\tfrac{i}{2}[\gamma_{\mu},\gamma_{\nu}]$ and $\sigma_{q\nu}=\sigma_{\mu\nu}q^{\mu}$)
@@
Lorentz structures ensuring translational invariance $q^{\mu}\Theta_{\mu\nu}(q)=0$. The normalisation $A(0)=1$ holds model-independently since $T_{\mu\nu}$ is the associated Noether current with conserved momenta $P_{\mu}=\int d^{3}xT_{\mu 0}$. Furthermore, for the nucleon one has $J^{N}(0)=\tfrac{1}{2}$, which reflects the absence of an anomalous gravitational magnetic moment, a consequence of the universality of gravity. [^3] The form factors $A^{\pi,N}$ and $J^{N}$ are analogues of the electromagnetic pion, Dirac and Pauli form factors.
@@ 2.2 Leading order dilaton effective theory and $D(q^{2})$
The form factors have been evaluated in the LO dilaton effective theory in an arbitrary spacetime dimension $d$, including chiral corrections, in ref. Zwicky:2023fay . Here, we consider it instructive to highlight the main effect, the dilaton pole in the $D$-form factor, using the nucleon as an example. Denoting the dilaton by $\sigma$ and the coset field by $\hat{\chi}\equiv e^{-\sigma/F_{\sigma}}$, the relevant Lagrangian reads
@@
where $F_{\sigma}$, $R$, $\omega_{N}$ and $\Delta_{\mu}$ stand for the dilaton decay constant, the Ricci scalar, the nucleon Weyl weight (conformal charge) and the Weyl-covariant derivative, respectively. For the latter we refer the reader to original refs. Mack:1969rr ; Isham:1970gz ; Isham:1970xz , and Zwicky:2023fay for an emphasis on generic Weyl-weight. We first focus on the improvement term proportional to the Ricci scalar. While irrelevant for scattering in flat space, it contributes to the energy-momentum tensor
@@
$D^{N}(q^{2})=\frac{r_{\sigma}^{N}}{q^{2}}\;,\qquad r_{\sigma}^{N}=\frac{2}{3}\bar{u}(p)u(p)\,F_{\sigma}\,g_{\sigma NN}=\frac{4}{3}m_{N}^{2}\;,$ (2.8)
@@
After illustrating the mechanism, let us turn to an LO-estimate, including linear quark mass effects. They have been determined in Zwicky:2023fay : [^4] Alternatively, one may write ${D}^{\pi}(q^{2})=\frac{2}{3}\frac{m_{\sigma}^{2}}{q^{2}-m_{\sigma}^{2}}-\frac{1}{3}$, which is more natural from the dispersive viewpoint.
$\quad{D}^{N}(q^{2})=\frac{4}{3}\frac{\bar{m}_{N}^{2}}{q^{2}-m_{\sigma}^{2}}\;,\quad{D}^{\pi}(q^{2})=\frac{2}{3}\frac{q^{2}}{q^{2}-m_{\sigma}^{2}}-1\;.$ (2.9)
The quantity $\bar{m}_{N}$ denotes the nucleon mass in the chiral limit
@@
and $\delta{m}_{N}$ the non-vanishing part due to $m_{u,d,s}\neq 0$. In the nucleon case the linear quark corrections vanish since the residue is of the form
$r_{\sigma}^{N}=\frac{4}{3}m_{N}(\bar{m}_{N}-\delta{m}_{N})+{\cal O}(m_{q}^{3/2})=\frac{4}{3}(\bar{m}_{N}^{2}-(\delta{m}_{N})^{2})+{\cal O}(m_{q}^{3/2})\;,$ (2.11)
which relies on the determination of the fixed-point anomalous dimension $\gamma_{*}=1$ of the quark mass Zwicky:2023bzk ; Zwicky:2023krx . The indicated corrections in (2.11) are radiative and of the same form as in baryon chiral perturbation theory Scherer:2012xha , whereas in the pion case they are of order ${\cal O}(m_{q}^{2}\ln m_{q})$ as usual.
Regarding the pion, it is worthwhile to point out that there is a soft-pion theorem constraining the LO trace to be Donoghue:1991qv ; Zwicky:2023fay
@@
Together with the model-independent normalisation $A^{\pi}(0)=1$, this implies the constraints
@@
subject to higher order chiral corrections. We emphasise that, for nonzero pion mass, the constraint holds independently of any fixed-point interpretation and is therefore directly relevant for the present work. In fact, Donoghue and Leutwyler Donoghue:1991qv employed the soft theorem together with the $\sigma$-meson dominance assumption to deduce this form. Consequently, the constraint $D^{\pi}(0)=-1$ provides a valuable handle for fixing the fit ansatz of the pion. In contrast, for the nucleon no analogous constraint exists, since constant or $q^{2}$-dependent shifts may arise from higher resonances or multi-particle contributions to the spectrum.
@@ 3 The $\sigma$-pole in the Euclidean
Describing the $\sigma$-pole is a non-trivial task since the $\sigma$-meson is perhaps the most complicated and mysterious resonance of QCD Pelaez:2015qba and its pole is deep in the complex plane on the second sheet Caprini:2005zr
@@
Our main point is that the details of the $\sigma$-meson in the Minkowski domain are irrelevant in the deep Euclidean region, where a simple effective pole parametrisation proves sufficient. In section 3.2, we will revisit the issue and explain why alternative resonance parametrisations are not well suited.
@@ 3.1 The effect of a resonance in the deep Euclidean
where $\cal O$} denotes the operator, transitioning from $H(p)$ to $H(p^{\prime})$, and coupling to a stable particle $\varphi$ in the $q^{2}=(p-p^{\prime})^{2}$ channel. We split the form factor into a part due to $\varphi$ and the rest $G(q^{2})=G_{\varphi}(q^{2})+\delta G(q^{2})$. Since $G$ (and also $G_{\varphi}$) are real for $q^{2}<m_{\varphi}^{2}$ (real analyticity), we may write the following dispersion representation
$G_{\varphi}(q^{2})=\int_{0}^{\infty}\frac{ds\,\rho_{\varphi}(s)}{s-q^{2}-i0}=\frac{r_{0,\varphi}}{q^{2}-m^{2}_{\varphi}}\;,$ (3.4)
@@
with spectral function $\rho_{\varphi}(s)=\tfrac{1}{\pi}\textrm{Im}G_{\varphi}(s)$. In this case, the effect of the pole across the complex plane is simple and fully controlled by the residue $r_{0,\varphi}$. Now suppose we turn on a parameter that renders the particle $\varphi$ unstable. The pole then moves to the second sheet, $\sqrt{s_{\varphi}}=m_{\varphi}-\frac{i}{2}\Gamma_{\varphi}$, and one may define a complex-valued residue $r_{\varphi}$ through
$G_{\varphi}^{(\text{II})}(q^{2})|_{q^{2}\approx s_{\varphi}}=\frac{r_{\varphi}}{q^{2}-s_{\varphi}}+{\cal O}(1)\;,$ (3.5)
where the superscript indicates the second-sheet continuation. This complex pole, while interesting in its own right, does not provide an effective description of $G_{\varphi}(q^{2})$ in the Euclidean domain. This is evident from the fact that $G_{\varphi}(q^{2})$ is real in this region. From the dispersion relation one then finds
$G_{\varphi}(q^{2})=\int_{0}^{\infty}\frac{ds,\rho_{\varphi}(s)}{s-q^{2}-i0}=\frac{r_{\mathrm{E},\varphi}}{q^{2}-m_{\mathrm{E},\varphi}^{2}}+\dots\;,$ (3.6)
where the dots denote terms that are suppressed for sufficiently Euclidean $q^{2}$ (see appendix B for further details where it is phrased in terms of a multipole-type expansion in momentum space). The effective residue is given by $r_{\mathrm{E},\varphi}=-\int ds,\rho_{\varphi}(s)$. This quantity represents an average over the real spectral density of the particle and therefore cannot coincide with the complex-valued residue.
Indeed, we anticipate
@@
$(i)\;\;\text{arg}(\sqrt{s_{\sigma}})\leftrightarrow\text{arg}(r_{\varphi})\;,\qquad(ii)\;\;r_{\mathrm{E},\varphi}\neq|r_{\varphi}|\;,\qquad{(iii)\;\;r_{\mathrm{E},\varphi}\approx r_{\text{LO}}\;,}$ (3.7)
namely that the phase of the pole correlates with the phase of the residue, that the Euclidean residue can differ significantly from the modulus of the complex residue, and that the effective residue is well approximated by the LO Lagrangian (provided the $\sigma$-meson acts as a pseudo dilaton). More generally, one would expect $m_{\mathrm{E},\varphi}\approx|s_{\varphi}|$ and that $|r_{\mathrm{E},\varphi}/r_{\varphi}|<1$, with this effect becoming more pronounced as the width-to-mass ratio increases. All of points (i), (ii) and (iii) are supported by the linear $\sigma$-model presented in appendix A, see in particular table 3.
Point (ii) is crucial, since for the nucleon the complex residue is known and differs substantially in its absolute value from the dilaton residue; see section 4.1 for a more detailed discussion. As point (iii) is central to our work, we would like to comment further. The LO result receives self-energy and vertex corrections, just as in the linear $\sigma$-model (A.3). The effect of the self-energy is absorbed into the effective mass, while the vertex corrections induce a $q^{2}$ dependence that is difficult to compute reliably within the effective theory. The parametric dependence of the effective theory will enter the error estimates of the dilaton predictions discussed later on. The leading corrections due to the quark mass are taken into account and missing corrections again enter as a parametric uncertainty estimate.